OpenStudy (anonymous):

HEEEELLLLPPPP Which statements are true? Choose all answers that are correct. A. –15 + (–10) > –4 + (–20) B. –10 + (–11) > 3 + (–26) C. 17 + (–26) > 12 + (–19) D. –18 + 31 = 24 + (–11)

OpenStudy (anonymous):

MANN

OpenStudy (anonymous):

d is one

OpenStudy (anonymous):

b and d

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Which statements are true? Choose all answers that are correct. A. –15 + (–10) > –4 + (–20) B. –10 + (–11) > 3 + (–26) C. 17 + (–26) > 12 + (–19) D. –18 + 31 = 24 + (–11)

OpenStudy (anonymous):

Which statements are true? Choose all answers that are correct. A. –15 + (–10) > –4 + (–20) B. –10 + (–11) > 3 + (–26) C. 17 + (–26) > 12 + (–19) D. –18 + 31 = 24 + (–11)

OpenStudy (anonymous):

your welcome